# Insights into the structure and kinematics of a Milky Way-like galaxy

### v2

@@ 1.1 The inner Galaxy structure
Numerical models have supported these findings, with portail2017dynamical estimating a pattern speed of $\Omega_{\mathrm{p}}=40$ km s-1 kpc-1. Other simulations, like those by sormani2022stellar, refine this estimate to $\Omega_{\mathrm{p}}\sim 37.5$ km s-1 kpc-1, in line with observations of the Milky Way’s boxy/peanut-shaped bulge structure (clark2019tracing; clarke2022; li2022). These studies collectively reinforce the characterization of the Milky Way’s bar as a key component of its overall structure.
@@
In terms of kinematics, gas in a non-axisymmetric potential such as that of a barred galaxy follows two different types of closed orbits near the Galactic Centre: $x_{1}$-type that are elongated orbits parallel to the bar major-axis; and $x_{2}$-type orbits, perpendicular to the bar major-axis (binney1991understanding). These orbits result from the bar dynamics and are dependent on the bulge size and mass (see e.g. bureau1999nature). Indeed, the existence, size and extent of these $x_{2}$ orbits are known to be influenced by the location of the Inner Lindblad Resonance (ILR) (e.g. Contopoulus1989; athan92a). Sparke1987 argues that this family of orbits is nearly empty in the absence of gas, but when gas is present, the bar drives it inwards and settles into $x_{2}$ orbits (binney1991understanding). The conventional understanding of bar orbital structures claims that the primary types of bar orbits are quasi-periodic or regular orbits, originating from stable $x_{1}$ and $x_{2}$ orbits (e.g. Contopoulus1980; Athanassoula1983). Abbot2017 examined the orbits primarily responsible for a boxy/peanut or X-shaped bulge shape (as found in the centre of the MW), and find that between $19-23\%$ of the bar’s mass is linked to the X-shape, significantly influenced by various bar orbit families including non-resonant box, ‘banana’, ‘fish/pretzel’, and ‘brezel’ orbits. No single family accounts for all observed features, but since Valluri2016 showed that box orbits are predominant in bars, and highly adaptable due to their unique frequencies, these are proposed as the most significant in shaping the X-bulge.
1.2 The Galactic disc and spiral structure
@@
Turning onto the Galactic disc, the nature and number of spiral arms in galaxies such as our MW, are subject of intense study and debate. The formation and persistence of these spiral structures, as reviewed by dobbs2014dawes, are influenced by a variety of dynamical processes. Key theories include the density wave theory, which proposes that spiral arms are density enhancements rotating through the disk, and gravitational instabilities contribute to the spontaneous emergence of spiral patterns. Additionally, tidal interactions with other galaxies can induce spiral structures, while the interplay between gas and stellar dynamics further shapes their appearance. And finally, N-body simulations suggest that there can be a more transient nature to the spiral structure of galaxies, where spiral arms are more dynamic, short-lived features organised by gravitational forces and recurrent patterns regulated by the heating and cooling processes within the disc (e.g. Sellwood1984; baba2015dynamics; sellwood2019spiral; pettitt2020spiral). For this type of spiral structure, spiral arms are not single, well-defined entities - they are constructed as a combination of superimposed arm-segments, complicating their detection and identification. This diversity in formation mechanisms leads to a variety of spiral arm structures observed across the galaxy population, from grand design spirals to flocculent and multi-armed galaxies, underscoring the complexity in determining a universal model for spiral arm formation, and consequently the role that spiral arms might have on star formation.
@@
The core of the debate about the number of spiral arms in the Milky Way lies in determining whether our Galaxy is characterised by a predominantly two-arm spiral structure, as some studies propose (drimmel2000evidence), or if it instead features a more complex configuration with four or more arms (e.g. churchwell2009).
@@
The discrepancies on the observed position and number of arms can be attributed to several factors. For one, different observational methods and wavelengths highlight various components of the arms, such as star-forming regions versus overall mass distribution. Optical observations often emphasise the locations of young, massive stars due to their brightness, which can suggest a simpler two-arm structure. However, optical data can capture a broad range of stellar populations (drimmel2001). In contrast, infrared observations, like those from the Spitzer GLIMPSE survey, are sensitive to star-forming regions and older and cooler stars, as they are particularly good at penetrating dust. These reveal a more complex structure with four or more arms, including several fainter arms and spurs (benjamin2005; churchwell2009). Radio observations of neutral hydrogen, which map the overall gas mass more comprehensively, also support a multi-armed configuration with four or more arms (levine2006; Hong2022).
@@
Observations of the molecular gas distribution in our Galaxy, also suggest the existence of multiple arms (e.g. dame2001milky), although their nature is still unclear. For instance, results from the observations of the first Galactic quadrant (e.g. ROMAN-DUVAL2010; rigby2016chimps; rigby2019) reveal enhancements in the surface density of molecular gas along some arms, and lower linewidths in the inter-arms, suggesting that molecular clouds are dynamically forming in spiral arms and may be disrupted in inter-arm spaces, which is more in line with grand-design type of pattern. That, however, is in contrast with the results from the SEDIGISM survey towards the 4th quadrant (e.g. duarte2021sedigism; colombo2022sedigism), where no significant contrast was found between the properties of molecular clouds in spiral arms and inter-arm regions, pointing towards a more transient nature of the arms. This disparity, however, could be partly due to the first quadrant being more affected by the bar dynamics and inducing stronger trends.
@@
It is worth noting, however, that most observational studies (except for those that use parallax measurements), have very high uncertainties in placing objects in their 3D position within the Galaxy. In particular for the gas, observations use the observed line-of-sight velocities to infer their kinematic distances (by assuming an idealised rotation curve, thus can be heavily distorted by any proper-motion of the gas). Hence, in order to better understand the link between the observed kinematics of the gas, and its inferred 3D structure, Ramon-fox2018 performed high-resolution Smoothed Particle Hydrodynamics (SPH) simulations of gas dynamics within spiral arms of a Milky Way-type model with a fixed 4-arm spiral potential (i.e. mimicking a density wave), without a bar. Their findings indicate that the spiral arms induce net radial motions of $\Delta\mathrm{V}_{R}$ $\sim-9$ km s-1 and azimuthal motions of $\Delta\mathrm{V}_{\Phi}\sim 6$ km s-1 slower than the rotation curve. Such dynamics lead to systematic errors in kinematic distance estimates by $\sim 1$ kpc. Consequently, observers mapping cloud positions based on these distances might encounter distortions and systematic offsets from the actual structures of up to $\pm 2$ kpc. These results are, of course, valid for this type of imposed potential that mimics a density wave, and this could become more complex if indeed the spiral arms are more dynamic in nature, and if the bar potential affects the dynamics of the disc further. Thus, validating simulation outputs against empirical data from observations becomes crucial for refining our understanding of Galactic structure.
Missions like Gaia and surveys such as APOGEE (e.g., katz2018gaia; Majewski2017; Minniti2010) have provided unprecedented precision in measurements of positions, velocities, and chemical compositions of stars across the Galaxy. Particularly, the Gaia mission has been instrumental in the study of the kinematic structure of the local neighbourhood. Gaia Data Release 3 (DR3, GaiaDR32021) has enhanced our view of the Milky Way by detailing the motions of stars within a few kpc of the Sun. khanna2023measuring used Gaia DR3 to map the velocities of stars, revealing intricate patterns of streaming motions in the Galactic disc. These observations provide empirical evidence for large-scale movements such as radial migration and vertical oscillations, where stars move perpendicular to the Galactic plane. Further analysis of Gaia data has also shown that these kinematic patterns vary significantly across different sectors of the Galaxy. Antoja2021 showed how Gaia DR3 data could be used to trace the influence of the Galactic bar on nearby star orbits, indicating that the bar’s gravitational pull significantly affects star velocities in its vicinity. This interaction results in asymmetries and kinematic waves, observable as ripples in the velocity field of stars, which propagate through the disc (Bland-Hawthorn2016). drimmel2023 further explored the disc of the Milky Way, and highlighted the asymmetries in the velocity fields and spatial distributions of stars, attributing these features to non-uniform mass distributions and ongoing dynamical processes influenced by the Galactic bar and spiral arms, being more evident in the outer regions of the Galaxy. The data from Gaia DR3+ reveal a more turbulent and dynamically active Milky Way than previously thought, challenging existing models to account for the observed complexity.
@@
A complementary avenue for studying MW-type galaxies involves fully cosmological simulations, which follow the assembly history, merger events, and large-scale environment of haloes similar to that of the MW (e.g. grand2017auriga; Wang2015; Buck2020; Schaye2015; Stinson2012; Serrato2023). These models are invaluable for understanding the statistical diversity of MW-type galaxies and for assessing how cosmological accretion and mergers shape galactic evolution. However, they are not suitable for the purpose of helping interpret the specific properties of the MW, as that requires a fine-tuned MW-analogue, which reproduces the detailed structural and kinematic properties of the MW itself. Furthermore, the high computational cost of tracking the evolution of a cosmological system makes it prohibitive to re-simulate the same system at substantially higher resolution or under varied physical prescriptions. For these reasons, and as discussed in DuranCamacho2024, we adopt a controlled isolated-galaxy framework, which allows the initial conditions to be tuned such that the resulting system resembles the present-day MW as closely as possible. This approach provides the flexibility and resolution needed to build a practical dynamical model of the MW, suitable for studying the origin of its present-day structures and kinematic patterns.
@@
In DuranCamacho2024, we explored a large suite of MW-type models with live gravitational potentials, which enables a dynamic and self-consistent evolution of the gas and stars, unlike models with imposed external fixed potentials (e.g dobbs2013exciting; ridley2017nuclear; smith2020cloud). In that work, we identified the model that was the closest reproduction of the MW structures and terminal velocities as seen in longitude-velocity space (lv space). In this follow-up study, we further test whether this model provides a robust dynamical description of the MW by benchmarking it against additional observational constraints and against studies modelling specific Galactic components (e.g. sormani2022stellar; portail2017dynamical; ridley2017nuclear). We then investigate the role of purely galaxy-driven kinematics in shaping the agglomeration, evolution, and disruption of spiral patterns in the model (thus prior to the inclusion of chemistry, star formation, and feedback, which will be subject of future work). We use these results to help deepen our understanding of the interplay between stellar and gaseous motions across the disc, bulge, and bar, and place these results in the context of recent Gaia DR3 findings and ongoing debates on the nature of MW spiral structure.
@@ 2 Model Overview
The simulation analysed in this study was introduced in DuranCamacho2024, as their Model 4. For conciseness, here we will only provide a brief description of the numerical setup, and readers are directed to DuranCamacho2024 for a more in-depth description of all the parameters. The model was generated using the arepo numerical code (springel2005cosmological), using a live potential of dark matter particles (following a spherical Hernquist profile), star particles (distributed in a disc with an exponential surface density profile, and a spherical stellar bulge following an Hernquist profile), and gas cells (distributed in a disc with an exponential surface density profile that follows that of the stars). These initial conditions were generated using the makenewdisc code (springel2005modelling). For the specific parameters that define each of these distributions see DuranCamacho2024. From a suite of 15 models with varying initial stellar mass distributions, DuranCamacho2024 identified Model 4 (at a time of 2.6 Gyr) as the one that most accurately matched the main observed structure of the Milky Way, as seen in lv space, using 12CO and H I data obtained from dame2001milky and bekhti2016hi4pi. That model has a total stellar mass of $4.25\times 10^{10}$ M (of which 10% is in the bulge), a dark matter mass of $8.97\times 10^{11}$ M, and a total gas mass in the disc of $8.6\times 10^{9}$ M.
@@ 3 The inner galaxy
Building upon the work from DuranCamacho2024, here we investigate in more detail the inner galaxy structure of our simulation, in order to test where and how it is able to capture some of the properties of the inner MW. This is an important benchmark required to inform potential other applications of this model as a MW analogue. In this section, we focus our analysis on the structure and kinematics of both stars and gas, and compare them with observationally inferred results.
@@ 3.1 Stellar distribution
Figure 2: Stellar surface density profiles along the $x$ (left), $y$ (centre), and $z$ (right) axes of our model (red-dotted line) compared to the analytic sormani2022stellar model (yellow-crossed line) including the overall best fit of an S22-type profile to our model (shown as a black solid line), and its various components represented by a distinct colour-coded dashed line: disc in dark blue, bar1 in pink, bar2 in green, and bar3 (i.e. long bar) in light blue.
@@
Here we explore the stellar density distribution in the inner galaxy of our numerical model (detailed in Section 2). In particular, we conduct a detailed examination of the stellar bar/bulge structure in our model, and compare these results against the characteristics of the stellar bar structure derived from observations of the inner Galaxy. For this comparison, we use the work of sormani2022stellar, which presents an analytical description of the Milky Way’s stellar bar, based on the N-body model from portail2017dynamical, developed to match observational data (e.g. wegg2013mapping; wegg2015structure). The final analytical model from S22 is made out of four distinct components: an inner bulge/bar or X-shape structure (bar1 and bar2), a long bar (bar3), and an axisymmetric disc. In Appendix A, we also include a further comparison with the stellar profiles of ridley2017nuclear, who use hydrodynamical simulations to model the dynamics of the Central Molecular Zone (CMZ) consistent with the observed gas flows within the inner Galaxy (including the $x_{2}$-type orbits). That model is able to reproduce the Inner Lindblad Resonances (ILR) and the barred galactic component, and hence also represents a good reference for the stellar profile in the Galactic Centre.
@@
The left panel of Figure 1 shows the projected stellar surface density maps of our model across the three principal axes, where the bar is oriented along the $x$ axis. The right panel of this figure shows the residual map between our model and S22’s analytical distribution. Figure 2 then shows 1D slices of the surface density of our model along the $x$, $y$ and $z$ axes (in red), alongside the profile from S22 (in yellow), and the best fit of an S22-type of profile to our model (in black). Details of this fitting can be found in Appendix A. While our model is almost always showing lower surface densities than the S22 analytic model [^2] Note that S22 is tailored to the central Galaxy, and as such, it should not be taken as an accurate representation of the Milky Way’s disc (which is the component that dominates their $x$ and $y$ profiles beyond $\sim 4-5$ kpc). We therefore restrain our comparison to the components that dominate the central region, related to the bar., these discrepancies are not severe, with similar shapes for the distributions in the inner $\sim 5$ kpc, suggesting that our model can be adequately described with the same type of components - as shown by the black lines on Fig. 2. Notably, the largest differences are observed in the $z$-direction, where our model diverges most significantly from S22, showing discrepancies of up to $0.3$ dex in surface density, which suggest the surface density in our model can be up to a factor 2 lower. On the other hand, both $x$ and $y$ slices show a good agreement with the analytic model, specially within the $\pm 2$ kpc region. The deviations in those directions are less pronounced, where the surface density of our model is only $\sim 10-15\%$ lower than S22. The most substantial discrepancies in the $x$-direction primarily arise from the bar2 component at $\sim 3$ kpc (see green curve on left panel of Fig. 2), whereas differences in the $y$-direction are predominantly due to the disc component in the $\sim 4-7$ kpc region (see blue curve on central panel of Fig. 2), although noting that this component is likely less accurate in S22.
@@
Overall, we find that our model can be fit by a boxy/peanut bar component (bar1,2 combined), with an approximate extension of $\sim 2$ kpc, and a long bar component (described by bar3) with a half-length extension of $\sim 3$ kpc, which aligns well with both analytical models and other observations (e.g. bissantz2002spiral; wegg2013mapping). Generally, our model presents a thinner distribution than S22 and R17 models, with lower surface density within the inner $\sim 5$ kpc, and particularly so in the $z$-direction, where we find discrepancies of up to $0.3$ dex. A lower central mass could potentially impact the kinematics of the inner galaxy, which we will look into in the next Section.

### v3

@@ 1 Introduction
Numerical models have supported these findings, with Portail et al. (2017) estimating a pattern speed of $\Omega_{\mathrm{p}}=40$ km s-1 kpc-1. Other simulations, like those by Sormani et al. (2022), refine this estimate to $\Omega_{\mathrm{p}}\sim 37.5$ km s-1 kpc-1, in line with observations of the Milky Way’s boxy/peanut-shaped bulge structure (Clark et al., 2019; Clarke and Gerhard, 2022; Li et al., 2022). These studies collectively reinforce the characterization of the Milky Way’s bar as a key component of its overall structure.
@@
In terms of kinematics, gas in a non-axisymmetric potential such as that of a barred galaxy follows two different types of closed orbits near the Galactic Centre: $x_{1}$-type that are elongated orbits parallel to the bar major-axis; and $x_{2}$-type orbits, perpendicular to the bar major-axis (Binney et al., 1991). These orbits result from the bar dynamics and are dependent on the bulge size and mass (see e.g. Bureau and Freeman, 1999). Indeed, the existence, size and extent of these $x_{2}$ orbits are known to be influenced by the location of the Inner Lindblad Resonance (ILR) (e.g. Contopoulos and Grosbol, 1989; Athanassoula, 1992). Sparke and Sellwood (1987) argues that this family of orbits is nearly empty in the absence of gas, but when gas is present, the bar drives it inwards and settles into $x_{2}$ orbits (Binney et al., 1991). The conventional understanding of bar orbital structures claims that the primary types of bar orbits are quasi-periodic or regular orbits, originating from stable $x_{1}$ and $x_{2}$ orbits (e.g. Contopoulos and Papayannopoulos, 1980; Athanassoula et al., 1983). Abbott et al. (2017) examined the orbits primarily responsible for a boxy/peanut or X-shaped bulge shape (as found in the centre of the MW), and find that between $19-23\%$ of the bar’s mass is linked to the X-shape, significantly influenced by various bar orbit families including non-resonant box, ‘banana’, ‘fish/pretzel’, and ‘brezel’ orbits. No single family accounts for all observed features, but since Valluri et al. (2016) showed that box orbits are predominant in bars, and highly adaptable due to their unique frequencies, these are proposed as the most significant in shaping the X-bulge.
@@
Turning onto the Galactic disc, the nature and number of spiral arms in galaxies such as our MW, are subject of intense study and debate. The formation and persistence of these spiral structures, as reviewed by Dobbs and Baba (2014), are influenced by a variety of dynamical processes. Key theories include the density wave theory, which proposes that spiral arms are density enhancements rotating through the disk, and gravitational instabilities contribute to the spontaneous emergence of spiral patterns. Additionally, tidal interactions with other galaxies can induce spiral structures, while the interplay between gas and stellar dynamics further shapes their appearance. And finally, N-body simulations suggest that there can be a more transient nature to the spiral structure of galaxies, where spiral arms are more dynamic, short-lived features organised by gravitational forces and recurrent patterns regulated by the heating and cooling processes within the disc (e.g. Sellwood and Carlberg, 1984; Baba et al., 2013; Sellwood and Carlberg, 2019; Pettitt et al., 2020). For this type of spiral structure, spiral arms are not single, well-defined entities - they are constructed as a combination of superimposed arm-segments, complicating their detection and identification. This diversity in formation mechanisms leads to a variety of spiral arm structures observed across the galaxy population, from grand design spirals to flocculent and multi-armed galaxies, underscoring the complexity in determining a universal model for spiral arm formation, and consequently the role that spiral arms might have on star formation.
@@
The core of the debate about the number of spiral arms in the Milky Way lies in determining whether our Galaxy is characterised by a predominantly two-arm spiral structure, as some studies propose (Drimmel, 2000), or if it instead features a more complex configuration with four or more arms (e.g. Churchwell et al., 2009).
@@
The discrepancies on the observed position and number of arms can be attributed to several factors. For one, different observational methods and wavelengths highlight various components of the arms, such as star-forming regions versus overall mass distribution. Optical observations often emphasise the locations of young, massive stars due to their brightness, which can suggest a simpler two-arm structure. However, optical data can capture a broad range of stellar populations (Drimmel and Spergel, 2001). In contrast, infrared observations, like those from the Spitzer GLIMPSE survey, are sensitive to star-forming regions and older and cooler stars, as they are particularly good at penetrating dust. These reveal a more complex structure with four or more arms, including several fainter arms and spurs (Benjamin et al., 2005; Churchwell et al., 2009). Radio observations of neutral hydrogen, which map the overall gas mass more comprehensively, also support a multi-armed configuration with four or more arms (Levine et al., 2006; Hong et al., 2022).
@@
Observations of the molecular gas distribution in our Galaxy, also suggest the existence of multiple arms (e.g. Dame et al., 2001), although their nature is still unclear. For instance, results from the observations of the first Galactic quadrant (e.g. Roman-Duval et al., 2010; Rigby et al., 2016, 2019) reveal enhancements in the surface density of molecular gas along some arms, and lower linewidths in the inter-arms, suggesting that molecular clouds are dynamically forming in spiral arms and may be disrupted in inter-arm spaces, which is more in line with grand-design type of pattern. That, however, is in contrast with the results from the SEDIGISM survey towards the 4th quadrant (e.g. Duarte-Cabral et al., 2021; Colombo et al., 2022), where no significant contrast was found between the properties of molecular clouds in spiral arms and inter-arm regions, pointing towards a more transient nature of the arms. This disparity, however, could be partly due to the first quadrant being more affected by the bar dynamics and inducing stronger trends.
@@
It is worth noting, however, that most observational studies (except for those that use parallax measurements), have very high uncertainties in placing objects in their 3D position within the Galaxy. In particular for the gas, observations use the observed line-of-sight velocities to infer their kinematic distances (by assuming an idealised rotation curve, thus can be heavily distorted by any proper-motion of the gas). Hence, in order to better understand the link between the observed kinematics of the gas, and its inferred 3D structure, Ramón-Fox and Bonnell (2018) performed high-resolution Smoothed Particle Hydrodynamics (SPH) simulations of gas dynamics within spiral arms of a Milky Way-type model with a fixed 4-arm spiral potential (i.e. mimicking a density wave), without a bar. Their findings indicate that the spiral arms induce net radial motions of $\Delta\mathrm{V}_{R}$ $\sim-9$ km s-1 and azimuthal motions of $\Delta\mathrm{V}_{\Phi}\sim 6$ km s-1 slower than the rotation curve. Such dynamics lead to systematic errors in kinematic distance estimates by $\sim 1$ kpc. Consequently, observers mapping cloud positions based on these distances might encounter distortions and systematic offsets from the actual structures of up to $\pm 2$ kpc. These results are, of course, valid for this type of imposed potential that mimics a density wave, and this could become more complex if indeed the spiral arms are more dynamic in nature, and if the bar potential affects the dynamics of the disc further. Thus, validating simulation outputs against empirical data from observations becomes crucial for refining our understanding of Galactic structure.
Missions like Gaia and surveys such as APOGEE (e.g., Gaia Collaboration et al., 2018, 2021a; Majewski et al., 2017; Bland-Hawthorn and Gerhard, 2016; Minniti et al., 2010) have provided unprecedented precision in measurements of positions, velocities, and chemical compositions of stars across the Galaxy. In particular, Gaia Data Release 3 (DR3, Gaia Collaboration et al., 2021b) has enabled detailed kinematic maps within a few kpc of the Sun; for example, Khanna et al. (2023) revealed intricate streaming-motion patterns in the Galactic disc that likely encode the imprint of non-axisymmetric structures. Gaia Collaboration et al. (2023) further explored the disc of the Milky Way, and highlighted the asymmetries in the velocity fields and spatial distributions of stars, attributing these features to non-uniform mass distributions and ongoing dynamical processes influenced by the Galactic bar and spiral arms, being more evident in the outer regions of the Galaxy. At the same time, Gaia has accelerated efforts to map the Milky Way spiral structure using young stellar tracers, although the inferred morphology remains uncertain and tracer-dependent due to extinction, selection effects, and our location within the disc. A recent overview of these developments in Galactic dynamics and their interpretation in the Gaia era is provided by Hunt and Vasiliev (2025). Complementary spiral-arm constraints come from radio maser parallaxes (e.g., Reid et al., 2019; Marshall et al., 2025) and standard-candle populations such as classical Cepheids. In particular, Drimmel et al. (2025) present a recent parametric spiral-arm model based on Cepheids with mid-infrared distance estimates, extending spiral mapping to distances where optical tracers are limited (see also Xu et al., 2023). Overall, the data from Gaia DR3+ reveal a more turbulent and dynamically active Milky Way than previously thought, challenging existing models to account for the observed complexity.
@@
A complementary avenue for studying MW-type galaxies involves fully cosmological simulations, which follow the assembly history, merger events, and large-scale environment of haloes similar to that of the MW (e.g. Grand et al., 2017; Wang et al., 2015; Buck et al., 2020; Schaye et al., 2015; Stinson et al., 2012; Sarrato-Alós et al., 2023). These models are invaluable for understanding the statistical diversity of MW-type galaxies and for assessing how cosmological accretion and mergers shape galactic evolution. However, they are not suitable for the purpose of helping interpret the specific properties of the MW, as that requires a fine-tuned MW-analogue, which reproduces the detailed structural and kinematic properties of the MW itself. Furthermore, the high computational cost of tracking the evolution of a cosmological system makes it prohibitive to re-simulate the same system at substantially higher resolution or under varied physical prescriptions. For these reasons, and as discussed in Durán-Camacho et al. (2024), we adopt a controlled isolated-galaxy framework, which allows the initial conditions to be tuned such that the resulting system resembles the present-day MW as closely as possible. This approach provides the flexibility and resolution needed to build a practical dynamical model of the MW, suitable for studying the origin of its present-day structures and kinematic patterns.
@@
In Durán-Camacho et al. (2024), we explored a large suite of MW-type models with live gravitational potentials, which enables a dynamic and self-consistent evolution of the gas and stars, unlike models with imposed external fixed potentials (e.g Dobbs and Pringle, 2013; Ridley et al., 2017; Smith et al., 2020). In that work, we identified the model that was the closest reproduction of the MW structures and terminal velocities as seen in longitude-velocity space (lv space). In this follow-up study, we further test whether this model provides a robust dynamical description of the MW by benchmarking it against additional observational constraints and against studies modelling specific Galactic components (e.g. Sormani et al., 2022; Portail et al., 2017; Ridley et al., 2017). We then investigate the role of purely galaxy-driven kinematics in shaping the agglomeration, evolution, and disruption of spiral patterns in the model (thus prior to the inclusion of chemistry, star formation, and feedback, which will be subject of future work). We use these results to help deepen our understanding of the interplay between stellar and gaseous motions across the disc, bulge, and bar, and place these results in the context of recent Gaia DR3 findings and ongoing debates on the nature of MW spiral structure.
@@ 2 Model overview
The simulation analysed in this study was introduced in Durán-Camacho et al. (2024), as their Model 4. For conciseness, here we will only provide a brief description of the numerical setup, and readers are directed to Durán-Camacho et al. (2024) for a more in-depth description of all the parameters. The model was generated using the arepo numerical code (Springel, 2005), using a live potential of dark matter particles (following a spherical Hernquist profile), star particles (distributed in a disc with an exponential surface density profile, and a spherical stellar bulge following an Hernquist profile), and gas cells (distributed in a disc with an exponential surface density profile that follows that of the stars). These initial conditions were generated using the makenewdisc code (Springel et al., 2005). For the specific parameters that define each of these distributions see Durán-Camacho et al. (2024). From a suite of 15 models with varying initial stellar mass distributions, Durán-Camacho et al. (2024) identified Model 4 (at a time of 2.6 Gyr) as the one that most accurately matched the main observed structure of the Milky Way, as seen in lv space, using 12CO and H I data obtained from Dame et al. (2001) and HI4PI Collaboration et al. (2016). That model has a total stellar mass of $4.25\times 10^{10}$ M (of which 10% is in the bulge), a dark matter mass of $8.97\times 10^{11}$ M, and a total gas mass in the disc of $8.6\times 10^{9}$ M.
@@ 3 The inner galaxy
Building upon the work from Durán-Camacho et al. (2024), here we investigate in more detail the inner galaxy structure of our simulation, in order to test where and how it is able to capture some of the properties of the inner MW. This is an important benchmark required to inform potential other applications of this model as a MW analogue. In this section, we focus our analysis on the structure and kinematics of both stars and gas, and compare them with observationally inferred results.
@@ 3.1 Stellar distribution
Figure 2: Stellar surface density profiles along the $x$ (left), $y$ (centre), and $z$ (right) axes of our model (red-dotted line) compared to the analytic Sormani et al. (2022) model (yellow-crossed line) including the overall best fit of an S22-type profile to our model (shown as a black solid line), and its various components represented by a distinct colour-coded dashed line: disc in dark blue, bar1 in pink, bar2 in green, and bar3 (i.e. long bar) in light blue.
@@
Here we explore the stellar density distribution in the inner galaxy of our numerical model (detailed in Section 2). In particular, we conduct a detailed examination of the stellar bar/bulge structure in our model, and compare these results against the characteristics of the stellar bar structure derived from observations of the inner Galaxy. For this comparison, we use the work of Sormani et al. (2022, hereafter S22), which presents an analytical description of the Milky Way’s stellar bar, based on the N-body model from Portail et al. (2017), developed to match observational data (e.g. Wegg and Gerhard, 2013; Wegg et al., 2015). The final analytical model from S22 is made out of four distinct components: an inner bulge/bar or X-shape structure (bar1 and bar2), a long bar (bar3), and an axisymmetric disc. In Appendix A, we also include a further comparison with the stellar profiles of Ridley et al. (2017, hereafter R17), who use hydrodynamical simulations to model the dynamics of the Central Molecular Zone (CMZ) consistent with the observed gas flows within the inner Galaxy (including the $x_{2}$-type orbits). That model is able to reproduce the Inner Lindblad Resonances (ILR) and the barred galactic component, and hence also represents a good reference for the stellar profile in the Galactic Centre.
@@
The left panel of Figure 1 shows the projected stellar surface density maps of our model in the $xy$, xzxz, and $yz$ projections, with the bar aligned with the $x$-axis. The right panel of this figure shows the residual map between our model and S22’s analytical distribution. Figure 2 then presents 1D surface-density cuts along the $x$, $y$, and $z$ axes for our model (red), the S22 analytic model (yellow), and the best-fitting S22-type decomposition to our model (black; see Appendix A). While our model is almost always showing lower surface densities than the S22 analytic model [^2] Note that S22 is tailored to the central Galaxy, and as such, it should not be taken as an accurate representation of the Milky Way’s disc (which is the component that dominates their $x$ and $y$ profiles beyond $\sim 4-5$ kpc). We therefore restrain our comparison to the components that dominate the central region, related to the bar., the differences are modest in the inner $\sim 5$ kpc: the overall shapes of the profiles match well, and the S22-type decomposition provides an adequate description of our model (black curves in Fig. 2). The largest discrepancy is in the vertical direction: the $z$ cut in Fig. 2 (third panel) differs by up to $\sim 0.3$ dex, i.e. our model can be up to a factor $\sim 2$ lower in surface density. In the plane, the agreement is better: within $|x|\lesssim 2$ kpc and $|y|\lesssim 2$ kpc (left and middle panels of Fig. 2), our model is typically only $\sim 10$–15% below S22. The dominant residual in the $x$ direction is associated with the bar2 component around $|x|\sim 3$ kpc (green curve; Fig. 2, left panel), whereas in the $y$ direction the main difference appears at $|y|\sim 4$–7 kpc where the disc component dominates in S22 (blue curve; Fig. 2, middle panel), noting that this disc component is less constrained in S22.
@@
Overall, we find that our model can be fit by a boxy/peanut bar component (bar1,2 combined), with an approximate extension of $\sim 2$ kpc, and a long bar component (described by bar3) with a half-length extension of $\sim 3$ kpc, which aligns well with both analytical models and other observations (e.g. Bissantz and Gerhard, 2002; Wegg and Gerhard, 2013). Generally, our model presents a thinner distribution than S22 and R17 models, with lower surface density within the inner $\sim 5$ kpc, and particularly so in the $z$-direction, where we find discrepancies of up to $0.3$ dex. A lower central mass could potentially impact the kinematics of the inner galaxy, which we will look into in the next Section.